The following ANOVA table is from a multiple regression analysis:

​​​​​​​MS F Source Regression Error Total df 5 25 SS 2000 2500 The observed F value is __ O 20 O 400 O 2000 O 500 O 10
Answer #1

To complete the given ANOVA table we use following formulae:

Total SS = Regression SS + Error SS

Hence, Error SS = Total SS - Regression SS = 2500 - 2000 = 500

MSS=\frac{SS}{df}

Regression MSS = 2000/5 = 400

Error SS = 500 / 25 = 20

F=\frac{RegressionMSS}{ErrorMSS}=\frac{400}{20}=20

ANOVA TABLE

Source df SS MSS F
Regression 5 2000 400 20
Error 25 500 20 -
Total 30 2500 - -

The observed F-value is 20.

I hope you find the solution helpful. If you have any doubt then feel free to ask in the comment section.

Please do not forget to vote the answer. Thank you in advance!!!